McCann Catching, Inc. has 2.00 million shares of stock outstanding. The stock currently sells for $12.94 per share. The firm’s debt is publicly traded and was recently quoted at 92.00% of face value. It has a total face value of $14.00 million, and it is currently priced to yield 10.00%. The risk free rate is 4.00% and the market risk premium is 8.00%. You’ve estimated that the firm has a beta of 1.18. The corporate tax rate is 35.00%.

The firm is considering a $40.71 million expansion of their production facility. The project has the same risk as the firm overall and will earn $11.00 million per year for 6.00 years.

What is the WACC for McCann Catching, Inc.?

What is the NPV of the expansion? (answer in terms of millions, so 1,000,000 would be 1.0000)
Answer #1

Common Stock: Number of Shares Outstanding = 2 million, Stock Price = $ 12.94, Value of Common Stocks = E = 12.94 x 2000000 = $ 25880000, RIsk-Free Rate = Rf = 4 %, Beta = 1.18, Market Risk Premium = MRP = 8 %

Cost of Common Stock = ke = Rf + Beta x MRP = 4 + 1.18 x 8 = 13.44 %

Debt: Face Value of Debt = $ 14 million, Market Value = D = 92 % of Face Value = 0.92 x 14000000 = $ 12880000

Cost of Debt = Yield = 10 % = kd

Firm Value = V = E + D = 25880000 + 12880000 = $ 38760000

Tax Rate = 35 % = t

WACC = kd x (1-t) x (D/V) + ke x (E/V) = 10 x (1-0.35) x (12880000/38760000) + 13.44 x (25880000/38760000) = 11.299979 % ~ 11.3 %

Expansion Project: Initial Investment = $ 40.71 million, Annual Cash Flows = $ 11 million and Cash Flow tenure = 6 years

Therefore, Project NPV = 11 x (1/0.1129) x [1-{1/(1.1129)^(6)}] - 40.71 = $ 5.43972 million ~ $ 5.44 million